# Benchmark Leakage Trap: Can We Trust LLM-based Recommendation?

### v1

@@ 1 Introduction
Recently, the Large Language Models (LLMs) (e.g., GPT-4 OpenAI (2023), LLaMA Touvron et al. (2023)) demonstrates unprecedented capabilities in natural language understanding, which has been employed in various fields. The recommender systems based on large language models (LLMs) have emerged as a promising direction Wu et al. (2023a); Wang et al. (2024). These models are pre-trained on massive corpora and encode rich knowledge information, which could enhance the recommender systems by deeper understanding of item content, user intent and contextual nuances Geng et al. (2022b); Li et al. (2023a); Liu et al. (2023); Geng et al. (2022a). For example, BinLLM Zhang et al. (2024) integrates collaborative information into LLMs using a text-like encoding strategy for recommendation, which has demonstrated superior performance.
However, should the results of these LLM-based recommender systems really be considered reliable? Recent studies have revealed that large language models are susceptible to data leakage, where models inadvertently memorize and reproduce training data during inference Carlini et al. (2021); Hisamoto et al. (2020); Carlini et al. (2022). This phenomenon has been extensively documented in various domains, from text generation Carlini et al. (2023) to question answering Lee et al. (2022), where LLMs have been shown to reproduce verbatim passages from their training corpora. And it largely affects the evaluation of model performance Zhou et al. (2023). Although there is ample evidence that LLMs are exposed to benchmark data during pre-training and affect downstream inference resultsCarlini et al. (2021); Lee et al. (2022), it remains unclear how leaked LLMs affect downstream recommendation results when used as the backbone model of recommendation systems.
@@ 2.1 Large Language Models
Large Language Models (LLMs) are a class of foundational models pre-trained on vast corpora of text data, demonstrating remarkable capabilities in understanding and generating human-like text OpenAI (2023); Bommasani et al. (2021); Brown et al. (2020). These models, typically based on the Transformer architecture, learn to predict the next token in a sequence, thereby capturing intricate patterns, facts, and reasoning abilities embedded in the training data Wei et al. (2022); Kaplan et al. (2020). A significant security and privacy concern associated with these models is data leakage, wherein LLMs can store and subsequently reproduce verbatim content from their training datasets Carlini et al. (2021). The study of data leakage in LLMs has become a critical area of research. Seminal work by Carlini et al.  Carlini et al. (2021) first systematically demonstrates that LLMs can memorize and regurgitate training data, showing that simple prompting strategies could recover exact sequences.  Zhou et al. (2023) shows that data leakage of large models will interfere with subsequent performance evaluation. And in response to address this problem, mitigation strategies like differential privacy and training data deduplication have been proposed to reduce these risks Lee et al. (2022).
@@ 2.2 LLM-based Recommendation
The field of LLM-based recommendation has evolved rapidly, with researchers exploring various architectural paradigms to integrate language models with recommendation tasks Wu et al. (2023b); Lin et al. (2023). LLMRec and LLMRec+Collab. are two primary architectural paradigms, which are directly represented in our experimental evaluation.
@@
In contrast, the second category, LLMRec+Collab. methods, explicitly integrates collaborative filtering signals to enhance the LLM’s reasoning with user-item interaction patterns. The integration strategies, however, vary significantly: PersonPrompt Li et al. (2023b) injects collaborative information by introducing trainable token embeddings for users and items, effectively performing personalized soft-prompt tuning. CoLLM Zhang et al. (2025), in its two instantiations, maps pre-trained collaborative embeddings from Matrix Factorization (CoLLM-MF) or a Deep Interest Network (CoLLM-DIN) into the LLM’s latent space. A distinctly different approach is taken by BinLLM Zhang et al. (2024), which encodes collaborative signals into a text-like format of binary hash codes for direct inclusion in the model’s prompt. This clear taxonomic distinction provides the essential context for interpreting our experimental results on how these fundamentally different paradigms respond to data leakage scenarios.
@@ 4.1.1 Datasets
  • Amazon-Book Ni et al. (2019b): This dataset is derived from the "Books" category of the well-known Amazon Product Review dataset. It comprises book purchasing and rating records, user-book interactions within the e-commerce domain.
For simulating the data leakage, we employ six external datasets from different domains to serve as out-of-domain leakage data, including Epinions (2010) (SNAP), Last.fm Celma (2010), MIND Wu et al. (2020), Amazon-Sports Ni et al. (2019a), Amazon-Beauty Ni et al. (2019a) and Gowalla (2011) (SNAP). The detailed information of the above datasets is presented in Table 1.
@@ 4.1.2 Baselines
  • PersonPrompt Li et al. (2023b): An LLMRec method that injects collaborative signals by introducing new tokens and embeddings for users/items. It can be viewed as personalized soft-prompt tuning on top of an LLM.
@@ 4.2.2 Ablation Study
  • Epinions: Built based on user history (2010) (SNAP) (the same as the LM-1M).
  • Last.fm: Built based on user interest tags Celma (2010).
  • Amazon-Sports: Built based on product attributes Ni et al. (2019a).

### v2

@@ 1 Introduction
Recently, the Large Language Models (LLMs) (e.g., GPT-4 Achiam et al. (2023), LLaMA Touvron et al. (2023)) demonstrates unprecedented capabilities in natural language understanding, which has been employed in various fields. The recommender systems based on large language models (LLMs) have emerged as a promising direction Wu et al. (2024); Wang et al. (2024). These models are pre-trained on massive corpora and encode rich knowledge information, which could enhance the recommender systems by deeper understanding of item content, user intent and contextual nuances Geng et al. (2022); Li et al. (2024); Liu et al. (2023). For example, BinLLM Zhang et al. (2024) integrates collaborative information into LLMs using a text-like encoding strategy for recommendation, which has demonstrated superior performance.
However, should the results of these LLM-based recommender systems really be considered reliable? Recent studies have revealed that large language models are susceptible to data leakage, where models inadvertently memorize and reproduce training data during inference Carlini et al. (2021); Hisamoto et al. (2020); Carlini et al. (2022). This phenomenon has been extensively documented in various domains, from text generation Carlini et al. (2023) to question answering Kandpal et al. (2022), where LLMs have been shown to reproduce verbatim passages from their training corpora. And it largely affects the evaluation of model performance Zhou et al. (2023). Although there is ample evidence that LLMs are exposed to benchmark data during pre-training and affect downstream inference resultsCarlini et al. (2021); Kandpal et al. (2022), it remains unclear how leaked LLMs affect downstream recommendation results when used as the backbone model of recommendation systems.
@@ 2.1 Large Language Models
Large Language Models (LLMs) are a class of foundational models pre-trained on vast corpora of text data, demonstrating remarkable capabilities in understanding and generating human-like text Achiam et al. (2023); Bommasani et al. (2021); Brown et al. (2020). These models, typically based on the Transformer architecture, learn to predict the next token in a sequence, thereby capturing intricate patterns, facts, and reasoning abilities embedded in the training data Wei et al. (2022); Kaplan et al. (2020). A significant security and privacy concern associated with these models is data leakage, wherein LLMs can store and subsequently reproduce verbatim content from their training datasets Carlini et al. (2021). The study of data leakage in LLMs has become a critical area of research. Seminal work by Carlini et al.  Carlini et al. (2021) first systematically demonstrates that LLMs can memorize and regurgitate training data, showing that simple prompting strategies could recover exact sequences.  Zhou et al. (2023) shows that data leakage of large models will interfere with subsequent performance evaluation. And in response to address this problem, mitigation strategies like differential privacy and training data deduplication have been proposed to reduce these risks Kandpal et al. (2022).
@@ 2.2 LLM-based Recommendation
The field of LLM-based recommendation has evolved rapidly, with researchers exploring various architectural paradigms to integrate language models with recommendation tasks Wu et al. (2024); Lin et al. (2025). LLMRec and LLMRec+Collab. are two primary architectural paradigms, which are directly represented in our experimental evaluation.
@@
In contrast, the second category, LLMRec+Collab. methods, explicitly integrates collaborative filtering signals to enhance the LLM’s reasoning with user-item interaction patterns. The integration strategies, however, vary significantly: Personalized recommendation Lyu et al. (2024) injects collaborative information by introducing trainable token embeddings for users and items, effectively performing personalized soft-prompt tuning. CoLLM Zhang et al. (2025), in its two instantiations, maps pre-trained collaborative embeddings from Matrix Factorization (CoLLM-MF) or a Deep Interest Network (CoLLM-DIN) into the LLM’s latent space. A distinctly different approach is taken by BinLLM Zhang et al. (2024), which encodes collaborative signals into a text-like format of binary hash codes for direct inclusion in the model’s prompt. This clear taxonomic distinction provides the essential context for interpreting our experimental results on how these fundamentally different paradigms respond to data leakage scenarios.
@@ 4.1.1 Datasets
  • Amazon-Book Ni et al. (2019): This dataset is derived from the "Books" category of the well-known Amazon Product Review dataset. It comprises book purchasing and rating records, user-book interactions within the e-commerce domain.
For simulating the data leakage, we employ six external datasets from different domains to serve as out-of-domain leakage data, including Epinions, Last.fm Celma Herrada and others (2009), MIND Wu et al. (2020), Amazon-Sports Ni et al. (2019), Amazon-Beauty Ni et al. (2019) and Gowalla. The detailed information of the above datasets is presented in Table 1.
@@ 4.1.2 Baselines
  • PersonPrompt Lyu et al. (2024): A LLMRec method that injects collaborative signals by introducing new tokens and embeddings for users/items. It can be viewed as personalized soft-prompt tuning on top of an LLM.
@@ 4.2.2 Ablation Study
  • Epinions: Built based on user history (the same as the LM-1M).
  • Last.fm: Built based on user interest tags Celma Herrada and others (2009).
  • Amazon-Sports: Built based on product attributes Ni et al. (2019).